OpenStudy (jojo12g):

What is the slope of the graph of 8x – 4y = 11? A. 2 B.-2 C.-1/2 D. 1/2

Mehek (mehek14):

Isolate y

OpenStudy (jojo12g):

i think the answer is D

OpenStudy (memyselfandpi):

Isolate y, then find what x is.

OpenStudy (memyselfandpi):

(What x is being multiplied by)

Mehek (mehek14):

\(8x – 4y = 11\\-4y=-8x+11\\ -y = -2x + \dfrac{11}{4} \\ y = 2x - \dfrac{11}{4}\)

OpenStudy (memyselfandpi):

Find what the value multiplying x is in mehek's equation.

OpenStudy (jojo12g):

so then its 2

OpenStudy (memyselfandpi):

Yes.

Mehek (mehek14):

Yep

OpenStudy (jojo12g):

thanks guys

Mehek (mehek14):

yw

OpenStudy (memyselfandpi):

Don't mention it.